OpenStudy (anonymous):

Can three segments of lengths 4cm,8cm, and 4square root 5 for a right triangle? Why or why not

OpenStudy (radar):

\[4\sqrt{5} \]Is that in cm also?

OpenStudy (radar):

Assuming that it is. \[\sqrt{4^{2}+8^{2}}=\sqrt{80}=\sqrt{16 \times5}=4\sqrt{5}\] Yup.

OpenStudy (anonymous):

yes it is.. thank you!